Find the derivative of the function f(x)=√x with respect to x at x0.

1
Expert's answer
2022-06-02T16:51:14-0400
"f'(x)=(\sqrt{x})'=\dfrac{1}{2\sqrt{x}}"

"f'(x_0)=\dfrac{1}{2\sqrt{x_0}}"